# Complex Magnetic Behavior in RuO2 Thin Films: Strain, Surface Effects, and Altermagnetic Signatures

### v1

@@ Abstract
RuO2 has been proposed as a prototypical metallic $d$-wave altermagnet, a Néel-ordered compensated antiferromagnetic state exhibiting nonrelativistic momentum-dependent spin splitting; yet, its magnetic ground state remains controversial both theoretically and experimentally. Using comprehensive first-principles calculations, we investigate RuO2 thin films with (110), (100), and (001) orientations, both freestanding and supported on a TiO2 substrate. We show that magnetism in RuO2 thin films is highly fragile, strongly influenced by strain, surface orientation, and atomic relaxation, while also being highly sensitive to the choice of the Brillouin-zone integration scheme. Although tensile strain induces finite magnetic moments, none of the studied systems stabilizes a compensated antiferromagnetic state; hence, an altermagnetic ground state cannot be achieved. Instead, the magnetic response is characterized by pronounced layer- and site-dependent variations and incomplete moment compensation, resembling a ferrimagnetic-like state. Our results reconcile conflicting theoretical and experimental reports and underscore the sensitivity of RuO2 magnetism to structural and methodological details.
@@ I Introduction
Finding and reliably identifying the true magnetic ground state of quantum materials remains one of the central challenges in condensed-matter physics and materials science [52, 4]. This task is particularly demanding in itinerant transition-metal oxides, where competing exchange interactions, electron itinerancy, and structural distortions often place the system close to multiple nearly degenerate magnetic instabilities [17, 49]. Even subtle perturbations, such as epitaxial strain, reduced dimensionality, surface termination, or details of electronic-structure methodology, can tip the balance between nonmagnetic, ferro-, ferri-, and antiferromagnetic states. Establishing the genuine magnetic ground state is, therefore, not only a fundamental problem but also a prerequisite for designing functional materials for spin-based technologies, where transport, symmetry, and magnetic order are intimately intertwined.
@@
Recently, antiferromagnetic (AFM) materials exhibiting nonrelativistic, momentum-dependent spin splitting have been identified as a symmetry-distinct class of antiferromagnets beyond conventional spin-degenerate AFM systems [36, 1, 35, 57, 44, 47, 45, 29, 3]. In these systems, the spin splitting originates not from spin–orbit coupling but from the spin space/point group symmetry, which allows for exchange-driven lifting of spin degeneracy while preserving zero net magnetization. The resulting momentum-dependent spin splitting is constrained by crystal symmetry and can be classified according to its angular-momentum character in the electronic band structure. In particular, unconventional collinear AFM materials with even-parity spin splitting of $d$-, $g$-, or $i$-wave symmetry, collectively termed altermagnets, retain fully compensated magnetic moments in real space while exhibiting sign-changing spin polarization across the Brillouin zone [47, 45]. This symmetry-enforced structure of the spin splitting distinguishes it from conventional $s$-wave AFMs, in which combined inversion or parity and time-reversal symmetries protect spin degeneracy throughout momentum space.
@@
Among these unconventional collinear AFMs, the $d$-wave class is particularly attractive because its symmetry allows for directional spin-polarized transport via the spin-splitter effect and spin splitting torque [16, 21]. Despite rapid progress in identifying candidate materials, many proposed altermagnets are insulating, which limits their immediate applicability in spin-transport devices. For practical applications, metallic or semiconducting systems are considerably more desirable, as they enable direct electrical control and detection of spin-polarized currents. RuO2 has been proposed as a prominent metallic candidate, with symmetry analysis predicting $d$-wave altermagnetic order. Early ab initio calculations [16, 46, 10] and some indirect quantum transport experimental observations [50, 10, 11, 18, 28, 60, 24, 23, 22] support this possibility. However, more recent high-resolution measurements, such as X-ray linear dichroism, nuclear magnetic resonance, spin-ARPES, and neutron diffraction; along with advanced first-principles studies, have challenged this picture, indicating that bulk RuO2 may, in fact, be nonmagnetic or, at most, extremely close to a magnetic [48, 40, 37, 20, 54]. This apparent contradiction has fueled an ongoing debate regarding the true magnetic ground state of RuO2 in bulk and thin films [8, 25, 33, 51, 27, 19, 55, 38, 31, 2, 7, 32, 30].
@@
Ab initio methods based on density functional theory (DFT) are powerful tools for determining magnetic ground states. Standard DFT calculations without an on-site Hubbard $U$ correction predict a nonmagnetic ground state for bulk RuO2, whereas DFT+$U$ calculations with $U$ values of $U>0.6$ eV stabilize an altermagnetic state [27, 42, 48]. Although the applicability of DFT+$U$, which effectively enhances the localization of Ru 4dd electrons, may be questioned for a system with predominantly itinerant character, the fact that magnetism emerges already for relatively small $U$ values indicates that RuO2 lies in close proximity to a magnetic instability. This pronounced sensitivity implies that its electronic and magnetic properties can be strongly influenced by external perturbations, including strain, symmetry breaking at surfaces and interfaces, and chemical doping [42, 12, 56, 20, 53].
@@
In this study, we perform comprehensive DFT calculations to achieve a microscopic understanding of the magnetic behavior of RuO2 thin films with (110), (001), and (100) orientations, both under strain and in the strain-free limit. Furthermore, we investigate RuO2/TiO2 thin-film heterostructures with the same crystallographic orientations in order to elucidate the role of epitaxial coupling and interface effects.
Since the (110) orientation is the most thermodynamically stable surface for both TiO2 and RuO2 [9, 58], our analysis places particular emphasis on this configuration. For this orientation, we systematically assess the influence of Brillouin-zone sampling, structural relaxation, induced magnetization, and competing magnetic ordering patterns in order to obtain a robust and quantitatively reliable determination of the magnetic ground state.
Our calculations demonstrate that RuO2 single crystals and thin films remain non-magnetic in the absence of external perturbations. Magnetic order may develop only within a few layers near the interface under tensile strain. Nevertheless, even under these conditions, the system does not realize a compensated altermagnetic phase, but instead adopts a ferrimagnetic-like state with a finite net magnetization.
II Method and Structures
II.1 Ab initio details
We performed nonrelativistic, extitab initio density functional theory (DFT) calculations using the Quantum ESPRESSO package [14, 13, 15]. The electron–ion interactions were described using the Standard Solid-State Pseudopotentials (SSSP) library, PBEsol Efficiency version 1.3.0 [41]. For the exchange–correlation functional, we employed the PBEsol approximation [39], a revised form of the Perdew–Burke–Ernzerhof (PBE) generalized gradient approximation (GGA) that provides improved accuracy for the structural properties of solids and surfaces. A vacuum spacing of approximately 15 Å was introduced in all thin-film calculations to avoid spurious interactions between periodic images. Brillouin-zone integrations were performed using a uniform $k$-point grid corresponding to a reciprocal-space resolution of 0.10 Å-1. The thin-film structures were constructed using the CIF2Cell code [5], and the initial Quantum ESPRESSO input files were generated with the QE input generator [41]. The plane-wave kinetic energy cutoff was set to 40 Ry for the wave functions and 400 Ry for the charge density.
We investigate three RuO2 thin-film surface orientations: (110), (100), and (001), considered both as freestanding slabs and supported on TiO2 substrates (see Fig. 1). The RuO2/TiO2 heterostructures are constructed using the experimental crystal structure of TiO2. For the RuO2(110)/TiO2(110) system, recently studied experimentally, we examine RuO2 film thicknesses ranging from one to seven Ru atomic layers. For the freestanding RuO2(110) thin film, a 17-layer Ru slab is employed to ensure convergence with respect to slab thickness and to minimize spurious interactions between the two surfaces. For the supported RuO2(100)/TiO2(100) and RuO2(001)/TiO2(001) heterostructures, the RuO2 film thickness is fixed at 12 Ru atomic layers. In the case of freestanding RuO2(100) and RuO2(001) thin films, thicker slabs consisting of 33 and 35 Ru atomic layers, respectively, are used to achieve well-converged surface properties.
@@ II.1 Ab initio details
For the RuO2(110) thin film, three distinct $(1\times 1)$ terminations of the rutile structure have been reported in the literature [43]. In this work, we consider two of these terminations. The relative stability of these terminations depends on the oxygen chemical potential, reflecting the oxygen availability during growth or operation [43]. Under oxygen-rich conditions, the stoichiometric RuO2(110)–Obridge termination is the most stable and is widely regarded as the preferred configuration of the rutile (110) surface. Under oxygen-poor conditions, the surface instead favors the RuO2(110)–Ocus termination, characterized by coordinatively unsaturated Ru sites (cus). The third possible termination, RuO2(110)–Ru, is not considered here, as it is unstable even under strongly oxygen-deficient conditions. The two terminations examined in this work are illustrated in Fig. 1.
Since our explicit comparison of oxygen-rich and oxygen-poor terminations for the (110) orientation reveals no significant changes in the magnetic ordering, we confine our analysis of the (100) and (001) thin films to terminations without oxygen atoms atop the Ru sites.
@@
Figure 1: Different surface orientations of RuO2 thin films considered in this work. For the RuO2(110)/TiO2(110) heterostructure, several film thicknesses were examined; the 5-L configuration is shown as a representative example. Blue and red spheres denote Ru atoms carrying different magnetic moments in the assumed AFM configurations discussed in the main text (labelled Ru1 and Ru2). Ti and O atoms are shown in green and white, respectively.
@@ III.1 Magnetism and k-point sampling sensitivity
A recent study [42] examined the influence of strain on the magnetic properties of bulk RuO2, demonstrating that lattice expansion stabilizes magnetic order—an effect that has also been reported in other materials [59]. Furthermore, Ref. [42] showed that the calculated magnetic moment in bulk RuO2 is highly sensitive to the choice of k-point sampling, exhibiting pronounced fluctuations that were attributed to the fragile and unconventional nature of magnetism in this compound.
Here, we show that a significant part of these fluctuations originates from the specific Brillouin-zone integration scheme employed. To substantiate this point, we compute the absolute magnetization of RuO2 using the lattice parameters and atomic positions of rutile TiO2. This choice effectively applies a tensile strain to RuO2, under which the system stabilizes in an AFM ground state.
We tested several Brillouin-zone integration schemes across different k-point meshes, including smearing-based approaches such as Marzari–Vanderbilt–DeVita–Payne cold smearing [34] and Fermi–Dirac smearing. Because smearing methods depend on the choice of the broadening parameter ($\sigma$), we also employed the optimized tetrahedron method [26], which enables accurate Brillouin-zone integration without smearing. The resulting absolute magnetizations were then systematically compared across all integration schemes and k-point grids.
As shown in Fig. 2, the calculated magnetization is highly sensitive to both the choice of the smearing scheme and the value of the broadening parameter $\sigma$. For example, Fermi–Dirac smearing with $\sigma=0.01$$0.03$ Ry yields a nearly vanishing magnetization, whereas cold smearing under the same conditions produces a large absolute magnetization of approximately 1.0 $\mu_{\mathrm{B}}$. These results underscore the importance of carefully selecting the Brillouin-zone integration scheme and broadening parameter when investigating magnetism in RuO2.
@@
Figure 2: Sensitivity of the Ru magnetic moment to k-point sampling in bulk rutile RuO2 under tensile strain. The calculations are performed for bulk RuO2 constructed using the lattice parameters and Wyckoff positions of rutile TiO2, which leads to a tensile strain and stabilizes an AFM state. k-point densities range from 0.05 Å-1 (corresponding to a $27\times 27\times 42$ Monkhorst–Pack mesh) to 0.30 Å-1 (corresponding to a $5\times 5\times 8$ mesh). (a) Ru magnetic moment as a function of the Monkhorst–Pack k-point grid calculated using Marzari–Vanderbilt–DeVita–Payne cold smearing for different smearing widths $\sigma$ (in Rydberg). Results obtained with the optimized tetrahedron method, which does not require a smearing parameter, are shown for comparison. (b) Same analysis performed using Fermi–Dirac smearing.
III.2 RuO2(110) thin films on TiO2(110)
Recent experimental investigations of the electronic band structure and transport properties of RuO2 have focused on RuO2(110) thin films grown on TiO2(110), owing to their favorable epitaxial matching and structural stability. We first investigate RuO2(110) thin films with thicknesses ranging from one to seven Ru-containing layers grown on a TiO2(110) substrate. Site-resolved magnetic moments are obtained by Bader partitioning of the spin density. In the AFM phase of bulk RuO2, the rutile structure contains two inequivalent Ru sites that form distinct magnetic sublattices, which we label Ru1 and Ru2. These sites are indicated in Fig. 1 by blue (Ru1) and red (Ru2) spheres.
Figure 3 presents the magnetic moments of Ru1 and Ru2 as a function of film thickness. In contrast to compensated AFM materials, the moments on Ru1 and Ru2 do not completely cancel. As a result, the RuO2(110) thin film grown on a TiO2(110) substrate develops a finite net magnetization, indicative of ferrimagnetic-like behavior.
Moreover, the magnetic moments exhibit a non-monotonic dependence on film thickness: the moments on both Ru1 and Ru2 oscillate as the number of Ru-containing layers increases, although without a strictly periodic pattern. These variations are most pronounced at the outermost surface layer and at the RuO2/TiO2(110) interface, underscoring the strong influence of surface and interface effects on the magnetic properties.

### v2

# Complex Magnetic Behavior in RuO2 Thin Films Driven by Strain and Substrate Effects

@@ Abstract
Ruthenium dioxide (RuO2) has been proposed as a prototypical metallic $d$-wave altermagnet, a Néel-ordered compensated antiferromagnetic state exhibiting nonrelativistic momentum-dependent spin splitting; yet, its magnetic ground state remains controversial both theoretically and experimentally. Using comprehensive first-principles calculations, we investigate RuO2 thin films with (110), (100), and (001) orientations, both (un)strained freestanding and supported on a TiO2 substrate. We show that emergent magnetic moments in RuO2 thin films are highly fragile, strongly influenced by strain, surface orientation, and atomic relaxation, while also being highly sensitive to the choice of the Brillouin-zone integration scheme. We find that none of the thin film structures considered can stabilize a compensated antiferromagnetic order; therefore, an altermagnetic ground state cannot be realized. Instead, substrate-supported RuO2 films on TiO2 exhibit pronounced layer- and site-dependent magnetic moment variations and incomplete compensation between the two antiferromagnetically coupled Ru moments, yielding a ferrimagnetic-like behavior. On the other hand, freestanding RuO2 films display complex magnetic structures depending on their orientation and applied strain, with distinct behavior at the surfaces and in the inner layers. Our results reconcile conflicting theoretical and experimental reports and underscore the sensitivity of RuO2 magnetism to structural and methodological details.
@@ I Introduction
Finding and reliably identifying the true magnetic ground state of quantum materials remains one of the central challenges in condensed-matter physics and materials science [56, 5, 4]. This task is particularly demanding in itinerant transition-metal oxides, where competing exchange interactions, electron itinerancy, and structural distortions often place the system close to multiple nearly degenerate magnetic instabilities [19, 53]. Even subtle perturbations, such as epitaxial strain, reduced dimensionality, surface termination, or details of electronic-structure methodology, can tip the balance between nonmagnetic, ferro-, ferri-, and antiferromagnetic states. Establishing the genuine magnetic ground state is, therefore, not only a fundamental problem but also a prerequisite for designing functional materials for spin-based technologies, where transport, symmetry, and magnetic order are intimately intertwined.
@@
Recently, unconventional antiferromagnetic (AFM) materials exhibiting nonrelativistic, momentum-dependent spin splitting have been identified as a symmetry-distinct class of antiferromagnets beyond conventional spin-degenerate AFM systems [38, 1, 37, 61, 48, 51, 49, 31, 4]. In these systems, the spin splitting originates not from the relativistic spin–orbit interaction but from the spin space/point group symmetry, which allows for exchange-driven lifting of spin degeneracy while preserving zero net magnetization. The resulting momentum-dependent spin splitting is constrained by crystal symmetry and can be classified according to its angular-momentum character in the electronic band structure. In particular, unconventional collinear AFM materials with even-parity spin splitting of $d$-, $g$-, or $i$-wave symmetry, collectively termed altermagnets, retain fully compensated magnetic moments in real space while exhibiting sign-changing spin polarization across the Brillouin zone [51, 49]. This symmetry-enforced structure of the spin splitting distinguishes it from conventional $s$-wave AFMs, in which combined inversion or parity and time-reversal symmetries protect spin degeneracy throughout momentum space.
@@
Among these unconventional collinear AFMs, the $d$-wave class is particularly attractive because its symmetry allows for directional spin-polarized transport via the spin-splitter effect and spin splitting torque [18, 23]. Despite rapid progress in identifying candidate materials, many proposed altermagnets are insulating, which limits their immediate applicability in spin-transport devices. For practical applications, metallic or semiconducting systems are considerably more desirable, as they enable direct electrical control and detection of spin-polarized currents. RuO2 has been proposed as a prominent metallic candidate, with symmetry analysis predicting $d$-wave altermagnetic order. Early ab initio calculations [6, 2, 50, 12] and a few direct experimental observations and indirect quantum transport measurements [6, 65, 54, 12, 13, 20, 30, 64, 26, 25, 24] support this possibility. However, more recent high-resolution measurements, such as X-ray linear dichroism, nuclear magnetic resonance, spin-ARPES, and neutron diffraction; along with advanced first-principles studies, have challenged this picture, indicating that bulk/thin-film RuO2 may, in fact, be nonmagnetic or, at most, extremely close to a magnetic [52, 44, 39, 22, 58]. This apparent contradiction has fueled an ongoing debate regarding the true magnetic ground state of RuO2 in bulk and thin films [10, 27, 35, 55, 29, 21, 59, 42, 33, 3, 9, 34, 32].
@@
Ab initio methods based on density functional theory (DFT) are powerful tools for determining magnetic ground states. Standard DFT calculations without an on-site Hubbard $U$ correction predict a nonmagnetic ground state for bulk RuO2, whereas DFT+$U$ calculations with $U$ values of $U>0.6$ eV stabilize a metallic altermagnetic state [29, 46, 52]. Although the applicability of DFT+$U$, which effectively enhances the localization of Ru 4dd electrons, may be questioned for a system with predominantly itinerant character, the fact that magnetism emerges already for relatively small $U$ values indicates that RuO2 lies in close proximity to a magnetic instability. This pronounced sensitivity implies that its electronic and magnetic properties can be strongly influenced by external perturbations, including strain, symmetry breaking at surfaces and interfaces, and chemical doping [46, 14, 60, 22, 57].
@@
In this study, we perform comprehensive DFT calculations to achieve a microscopic understanding of the magnetic behavior of both freestanding RuO2 slabs and epitaxial RuO2 thin films on TiO2 with (110), (001), and (100) orientations. Since the (110) orientation is the most thermodynamically stable surface for both TiO2 and RuO2 [11, 62], our analysis places particular emphasis on this configuration. For this orientation, we systematically assess the influence of Brillouin-zone sampling, structural relaxation, induced magnetic moments, and competing magnetic ordering patterns in order to obtain a robust and quantitatively reliable determination of the magnetic ground state.
Using systematically converged $k$-point sampling, we show that epitaxial RuO2 on a TiO2 substrate fails to stabilize a compensated altermagnetic state in all three orientations. Instead, it adopts a ferrimagnetic-like ground state characterized by a finite net magnetization and pronounced layer-dependent variations in the magnetic moment. In contrast, freestanding thin films remain nonmagnetic in the absence of strain, with only small magnetic moments emerging in a few surface-near layers; notably, the (001) surface remains entirely nonmagnetic. Under applied strain, freestanding films develop weak magnetic moments with varying amplitudes across the slab. Nevertheless, for the (001) orientation, the inner layers remain nonmagnetic, and only small surface-localized moments appear under strain.
II Model Structures
We investigate three RuO2 thin-film surface orientations, (110), (100), and (001); considered both as freestanding slabs and as substrate-supported films grown on TiO2 (see Fig. 1). The RuO2/TiO2 heterostructures are constructed using the experimental crystal structure of TiO2. For the RuO2(110)/TiO2(110) system, recently studied experimentally, we examine RuO2 film thicknesses ranging from one to seven Ru atomic layers. For the freestanding RuO2(110) thin film, a 17-layer Ru slab is employed to ensure convergence with respect to slab thickness and to minimize spurious interactions between the two surfaces. For RuO2(100)/TiO2(100) and RuO2(001)/TiO2(001) heterostructures, the RuO2 film thickness is fixed at 12 Ru atomic layers. In the case of freestanding RuO2(100) and RuO2(001) thin films, thicker slabs consisting of 33 and 35 Ru atomic layers, respectively, are used to achieve well-converged surface properties.
@@ II Model Structures
For the RuO2(110) thin film, three distinct $(1\times 1)$ terminations of the rutile structure have been reported in the literature [47]. In this work, we consider two of these terminations; see Fig. 1a. The relative stability of these terminations depends on the oxygen chemical potential, reflecting the oxygen availability during growth or operation [47]. Under oxygen-rich conditions, the stoichiometric RuO2(110)–Obridge termination is the most stable and is widely regarded as the preferred configuration of the rutile (110) surface. Under oxygen-poor conditions, the surface instead favors the RuO2(110)–Ocus termination, characterized by coordinatively unsaturated Ru sites (cus). The third possible termination, RuO2(110)–Ru, is not considered here, as it is unstable even under strongly oxygen-deficient conditions.
Since our explicit comparison of oxygen-rich and oxygen-poor terminations for the (110) orientation reveals no significant changes in the magnetic ordering, we limit our analysis of the (100) and (001) surfaces to terminations without oxygen atoms atop the Ru sites; see Figs. 1b and Fig. 1c.
@@
Figure 1: Different surface orientations of RuO2 thin films considered in this work, for both freestanding and substrate-supported configurations. For the RuO2(110)/TiO2(110) heterostructure, several film thicknesses were examined; the 5-L configuration is shown as a representative example. Blue and red spheres denote Ru atoms carrying different magnetic moments in the assumed AFM configurations discussed in the main text (labelled Ru1 and Ru2). Ti and O atoms are shown in green and white, respectively.
@@ III.1 Magnetism and k-point sampling sensitivity
A recent ab initio study [46] investigated the effect of strain on the magnetic properties of bulk RuO2, showing that lattice expansion stabilizes magnetic order, an effect also observed in related systems [63]. Moreover, the authors reported that the calculated magnetic moment is highly sensitive to the choice of k-point mesh, displaying pronounced fluctuations. This sensitivity was attributed to the fragile and unconventional character of magnetism in RuO2.
Here, we demonstrate that these fluctuations largely originate from the specific Brillouin-zone integration scheme employed. To substantiate this conclusion, we evaluate the total absolute magnetization of RuO2 under an effectively tensile-strained configuration obtained by adopting the lattice parameters and atomic positions of rutile TiO2. Here, the total absolute magnetization is defined as the sum over the absolute magnetic moments of all atoms in the unit cell, $|M|=\sum^{N_{\rm{atoms}}}_{i=1}|M^{i}|=\int_{\mathrm{cell}}\left|n_{\mathrm{up}}(\mathbf{r})-n_{\mathrm{down}}(\mathbf{r})\right|\,d^{3}r$, where $n_{\mathrm{up}}(\mathbf{r})$ and $n_{\mathrm{down}}(\mathbf{r})$ are the spin-up and spin-down charge densities. In this regime, RuO2 stabilizes a compensated AFM order and consequently exhibits altermagnetic behavior.
We tested several Brillouin-zone integration schemes across different k-point meshes, including smearing-based approaches such as Marzari–Vanderbilt–DeVita–Payne cold smearing [36] and Fermi–Dirac smearing. Because smearing methods depend on the choice of the broadening parameter ($\sigma$), we also employed the optimized tetrahedron method [28], which enables accurate Brillouin-zone integration without smearing. The resulting total absolute magnetizations were then systematically compared across all integration schemes and k-point grids.
As shown in Fig. 2, the calculated total absolute magnetization (and, consequently, the magnetic moments of the Ru atoms) is highly sensitive to both the choice of the smearing scheme and the value of the broadening parameter $\sigma$. For example, Fermi–Dirac smearing with $\sigma=0.01$$0.03$ Ry yields nearly vanishing magnetic moments, whereas cold smearing under the same conditions produces a large total absolute magnetization of approximately 1.0 $\mu_{\mathrm{B}}$. These results underscore the importance of carefully selecting the Brillouin-zone integration scheme and broadening parameter when investigating magnetism in RuO2.
@@
Figure 2: Sensitivity of the total absolute magnetization (and, consequently, the Ru magnetic moments) to k-point sampling in bulk rutile RuO2 under tensile strain. The calculations are performed for bulk RuO2 constructed using the lattice parameters and Wyckoff positions of rutile TiO2, which leads to a tensile strain and stabilizes an AFM state. k-point densities range from 0.05 Å-1 (corresponding to a $27\times 27\times 42$ Monkhorst–Pack mesh) to 0.30 Å-1 (corresponding to a $5\times 5\times 8$ mesh). (a) The total absolute magnetization as a function of the Monkhorst–Pack k-point grid calculated using Marzari–Vanderbilt–DeVita–Payne cold smearing for different smearing widths $\sigma$ (in Rydberg). Results obtained with the optimized tetrahedron method, which does not require a smearing parameter, are shown for comparison. (b) Same analysis performed using Fermi–Dirac smearing with a smearing amplitude $\sigma$.
III.2 RuO2(110)/TiO2(110) heterostructure
Recent experimental investigations of the electronic band structure and transport properties of RuO2 have focused on RuO2(110) thin films grown on TiO2(110), owing to their favorable epitaxial matching and structural stability. We first, therefore, investigate RuO2(110) thin films with thicknesses ranging from one to seven Ru-containing layers grown on a TiO2(110) substrate. Site-resolved magnetic moments are obtained by Bader partitioning of the spin density. In the AFM phase of bulk RuO2, the rutile structure contains two inequivalent Ru sites that form distinct magnetic sublattices, which we label Ru1 and Ru2. These sites are indicated in Fig. 1 by blue (Ru1) and red (Ru2) spheres.
Figure 3 presents the magnetic moments of Ru1 and Ru2 at each layer as a function of the film thickness for Bridge (top panel) and Cus (bottom panel) terminations. The opposite magnetic moments on Ru1 and Ru2, in both cases, do not completely cancel. Consequently, the RuO2(110) thin film on a TiO2(110) substrate exhibits a finite net magnetization, consistent with ferrimagnetic-like behavior rather than the compensated AFM order characteristic of altermagnets. We call it ferrimagnetic-like state since, contrary to a ferrimagnetic state, the magnetic moments exhibit layer-dependent amplitudes: the moments on both Ru1 and Ru2 in different layers oscillate as the thickness increases, although without a strictly periodic pattern. These variations are most pronounced at the outermost surface layer and at the RuO2/TiO2(110) interface, underscoring the strong influence of surface and interface effects on the magnetic properties.